OpenStudy (anonymous):

Change the given equation into slope-intercept form. 5x + 2y = 10

OpenStudy (anonymous):

5x + 2y = 10 slope intercept form in y = mx + b 2y = -5x + 10 Isolate y. 2y/2 = -5x/2 + 10/2 y = -5x/2 + 5 The answer is y = -5x/2 + 5.

OpenStudy (anonymous):

that isnt right thats not an option

OpenStudy (aravindg):

*y=mx +c

OpenStudy (aravindg):

rearrnge to get to it

OpenStudy (anonymous):

options are -5/2 5/2 2/5 -2/5

OpenStudy (anonymous):

It's asking for the value of the slope. The answer is -5/2.

OpenStudy (anonymous):

i was just bout to say that.. you beat me too it

OpenStudy (anonymous):

can u guys help me with one i just posted